A 1220-kg car pulls a 390-kg trailer. The car exerts a horizontal force of 3.6 × 103 N against the ground in order to accelerate.

What force does the car exert on the trailer? Assume an effective friction coefficient of 0.19 for the trailer.
Answer #1

writing the force equation we have

F - u*1200*g - T = 1200a -------(1)

T - u*390*g = 390a ---------(2)

adding both we have

3600 - 0.19*1590*9.8 = 1590a

a = 0.507 m/s2

T = 0.19*390*9.8 + 390*0.507

T = 923.9 N

so the force that the car exerts on the trailer is 923.9 N